OpenStudy (anonymous):

Please help! t(r)= d/r t=time in hours, d=distance, r=rate in mi/h Sydney drives 10 mi at a certain rate and then drives 20 mi at a rate 5 mi/h faster than the initial rate. Write expressions for the time along each part of the trip. Add these times to write an equation for the total time in terms of initial rate, t Total (r).

OpenStudy (anonymous):

\[\frac{10}{r}+\frac{20}{r+5}=\frac{30 r+50}{r^2+5 r} \]

OpenStudy (anonymous):

r+5 is the rate for when she drives 20 miles.

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Okay so what you put is the final answer correct?

OpenStudy (anonymous):

Yes, either side of the equation answers the question. The RHS is the result of combining the two fractions.

OpenStudy (anonymous):

Oh! Okay, good that is the same answer I got on my own, I just wanted to check it. Could you help me with the next part by any chance? It asks "At what rate, to the nearest mi/h, must Sydney drive for the first 10 miles if the entire 30 miles must be covered in about 45 minutes?

OpenStudy (anonymous):

45 minutes is 3/4 ths of an hour. Solve the following for r:\[\frac{10}{r}+\frac{20}{r+5}=\frac{3}{4} \]\[r=\frac{5}{6} \left(21+\sqrt{537}\right)=37 \text{ mph} \text{ rounded} \]

OpenStudy (anonymous):

Thank you so much! Lastly, "How long will it take Sydney to drive the entire 30 miles if the car's initial rate varies between 10 mi/h and 20 mi/h?

OpenStudy (anonymous):

The average speed is 30/2 or 15 mph.\[\frac{10 (3 r+5)}{r (r+5)} \]evaluated at r = 15 is 5/3 hour or 1 hour and 40 minutes

OpenStudy (anonymous):

At a constant speed of 10 mph and 20 mph, the total trip time would be 7/3 and 13/10 hours respectively.

OpenStudy (anonymous):

Thank you so much for your help!

OpenStudy (anonymous):

You're welcome.